# Entanglement dynamics and eigenstate correlations in strongly disordered quantum many-body systems

### v1

@@ I Introduction
The question of if, and how, out-of-equilibrium quantum many-body systems thermalise as quantum correlations develop dynamically between distant degrees of freedom is one of the most fundamental questions of interest in condensed matter and statistical physics [1, 2, 3, 4, 5, 6]. This question is intimately connected to the entanglement structure and its dynamics [7, 8, 9] as it is one of the most fundamental manifestations of how quantum correlations and information propagates through the system [10, 11, 12, 13, 14, 15, 16]. Developing a microscopic theory for the dynamics of entanglement in quantum many-body systems is therefore naturally important.
@@ III Disordered Floquet spin-1/2 chain as a model
While the entire framework relating the dynamics of purity to eigenstate correlations described in the previous section was completely general, our main focus on strongly disordered systems in the MBL regime. In order to orient ourselves in that direction, and also as a concrete setting for demonstrating our theory, we we employ a disordered, Floquet spin-1/2 chain. The Floquet unitary, $U_{F}$, is given by [65]
@@ IV.1 Special quartets which dominate dynamics of purity
The picture that therefore emerges for the structure of quartets of eigenstates with $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$ is as follows. The four eigenstates in the quartet are such that there localisation centres are made of two distinct $\sigma^{z}$-configurations in subsystem $A$ and two in $B$. In other words, the four distinct combinations of these states form the localisation centres of the four eigenstates. In fact, this picture straightforwardly explains the number of quartets with $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$. For a given |$\ket{\alpha}$, the number of states |$\ket{\beta}$ which has the same localisation centre in $B$ but different in $A$ is $N_{{\cal H}_{A}}-1$. Similarly, the number of states |$\ket{\gamma}$ which has the same localisation centre as |$\ket{\alpha}$ in $A$ but different in $B$ is $N_{{\cal H}_{B}}-1$. Fixing |$\ket{\beta}$ and |$\ket{\gamma}$ automatically fixes the state |$\ket{\lambda}$, see Eq. 28 or Fig. 5. Since there are $N_{\mathcal{H}}$ choices of |$\ket{\alpha}$ itself, the total number of quartets which satisfy Eq. 28, which we denote by ${\cal N}_{\square}$ is therefore $N_{\mathcal{H}}(N_{{\cal H}_{A}}-1)(N_{{\cal H}_{B}}-1)$. As the subsystems $A$ and $B$ correspond to left and right halves of the entire system, we have ${\cal N}_{\square}=N_{\mathcal{H}}(\sqrt{N_{\mathcal{H}}}-1)^{2}\approx N_{\mathcal{H}}^{2}$ or equivalently ${\cal N}_{\square}/{\cal N}_{Q}\sim N_{\mathcal{H}}^{-2}$. This immediately explains why a fraction $\sim N_{\mathcal{H}}^{-2}$ of the quartets have $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$ (see Fig. 4). It is worth clarifying and reiterating a couple of points here. First, our focus is deep in the MBL phase where associating an eigenstate to a localisation centre is well-defined despite the multifractality of the eigenstate. Second, the theoretical picture that emerges in the following sections does not depend on an eigenstate being sharply localised in the Hilbert space. Instead, the identification should be thought of as a starting point for constructing the theory.
@@ IV.3 Emergence of logarithmic growth of entanglement
While the summations in Eq. 42 can be done exactly, as detailed in Appendix C, here we only sketch the derivation and discuss the physical import of the result. A couple of key assertions that we make here are $\mu>1$ and $\xi<1/\ln 2$; the numerical results in Fig. 10 indeed show that they are satisfied. However, we will see post facto that violating these bounds leads to results which are unphysical or incompatible with locality.
In the first summation in Eq. 42, note that the summand, $\sim 2^{-r}e^{r/\xi}$ is an exponentially increasing function of $r$ as $\xi<1/\ln 2$, and hence the sum is dominated by $r=r_{\ast}(\omega)$. As such, the first sum can be estimated as $2^{-r_{\ast}(\omega)}e^{r_{\ast}(\omega)/\xi}r_{\ast}(\omega)$ which using Eq. 41 is $\sim\omega^{-1+\xi\ln 2}\ln|\omega|$. On the other hand, in the second summation, by virtue of $\mu>1$ and $\xi<1/\ln 2$, the summands are an exponentially decreasing function of $r$, and hence the summation is expected to be dominated by the first term $r=r_{\ast}(\omega)+1$, which again leads to an estimate of the sum of as $\sim\omega^{-1+\xi\ln 2}\ln|\omega|$, which is the same as the first summation. We therefore conclude that
$\langle\tilde{\mathcal{P}}^{AB}(\omega)\rangle\sim\omega^{-1+\xi\ln 2}|\ln\omega|\,,$ (43)
@@
$\langle\mathcal{P}^{AB}(t)\rangle\sim t^{-\xi\ln 2}\ln t\,,$ (44)
@@
$\langle S_{2}^{AB}(t)\rangle\sim\xi\ln 2\ln t-\ln\ln t\overset{t\gg 1}{\sim}(\xi\ln 2)\ln t\,,$ (45)
which is the logarithmic growth of the entanglement entropy characteristic of the MBL regime. From the data in Fig. 10, we have $\xi=0.32$ for $\Gamma=0.1$ and $\xi=0.44$ for $\Gamma=0.15$. For these values, the power-law-exponent, $\xi\ln 2$ in Eq. 44, takes on values approximately 0.22 and 0.31 for $\Gamma=0.1$ and $0.15$ respectively. The numerical results in Fig. 2 on the other hand estimate the exponent $a$ to be 0.17 and 0.24. We note that while the two estimates are reasonably close, $a\lessapprox\xi\ln 2$ which we attribute to the logarithmic correction in Eq. 44.
@@ IV.4 Spacetime picture of entanglement growth
An important physical import of the above analysis is that it provides a spacetime picture of the entanglement growth as follows. For a given $\omega$, the sum in Eq. 39 is dominated by a single $r=r_{\ast}(\omega)$. In other words, the dynamics of the purity at frequency $\omega$ is encoded in the spectral correlations of the form in Eq. 30 within the special quartets $\square_{r_{\ast}(\omega)}$ with $r_{\ast}(\omega)$ given by Eq. 41. This suggests that that at any time $t$, there is a a characteristic lengthscale $d(t)\equiv r_{\ast}(\omega=2\pi/t)\sim\xi\ln t$ and degrees of freedom within this distance $d(t)$ from the bipartition are entangled at time $t$. Since the total entanglement carried in such a case will be proportional to $d(t)$, this presents an alternative perspective to infer the logarithmic growth of entanglement entropy. In addition, this could also be understood as the manifestation of the logarithmic light-cone of operator spreading in the MBL regime [42, 43, 44, 45].
@@
where ${\cal P}_{r}(t)$ is the Fourier transfrom of $\tilde{\mathcal{P}}_{\square_{r}}(\omega)$. The functional form of $\tilde{\mathcal{P}}_{\square_{r}}(\omega)$ in Eq. 40 suggests that ${\cal P}_{r}(t)$ decays like a stretched-exponential in time [69],
@@
Since $t_{\ast}(r)$ grows exponentially with $r$, the temporal dependence of ${\cal P}_{r}(t)$ naturally slows down with increasing $r$. At the same time, the number of quartets with $r$ also decays exponentially with $r$, Eq. 35. The power-law decay of $\langle\mathcal{P}^{AB}(t)\rangle$ emerges from the collective effect of the stretched-exponential decays in Eq. 47 with hierarchically growing timescales, $t_{\ast}(r)\sim e^{r/\xi}$ and the hierarchically decreasing $N_{r}\sim e^{-r\ln 2}$. This arises simply from the fact that the sum in Eq. 46 is dominated by a ‘saddle-point’, $r_{\ast}(t)\sim\xi\ln t$. It is also interesting to note that the phenomenology is quite similar to that of hierarchically constrained dynamics for glassy relaxation [70].
The space-resolved picture discussed above also provides us with a notion of an ‘entanglement wavefront’. This can be understood as follows. The partial sum of Eq. 46, $\sum_{r=2}^{R}N_{r}{\cal P}_{r}(t)$ can be understood to contain the information of the decay of bipartite purity between $A$ and $B$ carried by degrees of freedom within a distance $2R$ from the bipartition. $N_{r}{\cal P}_{r}(t)$ is therefore the contribution to the decay in purity at time $t$ at distance $r$ from boundary. However, the aforementioned $r_{\ast}(t)$ suggests that there is a well-defined wavefront of entanglement spreading. This is shown schematically in Fig. 11 where $N_{r}{\cal P}_{r}(t)$ is plotted as a heatmap in spacetime assuming the stretched exponential in Eq. 47 which also shows the logarithmic spreading in time of the entanglement wavefront. For a given $t$, the wavefront decays exponentially with $r$ away from $r_{\ast}(t)$ for $r>r_{\ast}(t)$ due to the exponential decay with $r$ of $N_{r}$. On the other hand, the hole in the profile for $r$ away from $r_{\ast}(t)$ with $r<r_{\ast}(t)$ is purely due to the fact that degrees of freedom within distance $r$ of the bipartition are strongly entangled and the effective contribution to the purity from them has decayed to extremely small values.
Figure 11: The spacetime picture of entanglement growth. The left panel shows a heatmap in spacetime of $N_{r}{\cal P}_{r}(t)$, defined in Eq. 46. The wavefront spreads logarithmically in time as indicated by the red dashed line. The right panel shows wavefronts in space at fixed times (darker colours show later times) which are horizontal cuts of the figure in the left panel.
@@ VI Summary and Outlook
All the results summarised above are rooted in an understanding of the eigenstate and spectral correlations involving four different eigenstates. Questions to follow up in the future therefore arise where such correlations or their generalisations appear naturally. While we focused here on the entanglement entropy of a system with no conservation laws, a similar theory can be constructed for the unresolved issue of the dynamics number entropy in MBL systems with a conserved charge [71, 72, 73, 74]. In order to better understand the spatiotemporal structure of statistics of eigenstates, analogous to Ref. [37], one can study the correlation $V_{\alpha\beta\gamma\lambda}(V_{\alpha\beta\gamma\lambda}^{\prime})^{\ast}$ resolved in $\theta_{\alpha\beta\gamma\lambda}$ where $V_{\alpha\beta\gamma\lambda}^{\prime}$ is the same eigenstate correlation as defined in Eq. 8 but with a different bipartition of the system. In a similar spirit, one can introduce matrix elements of local operators and recast the OTOC or temporal fluctuations of local observables [75] from the point-of-view of eigenstate correlations. In the same direction of spatiotemporal structure of entanglement dynamics, it will be important to understand the origins of the putative stretched-exponential decay in time of purity resolved in $r$ (see Eq. 47). Usually such anomalous decays appear due to the presence of a broad distribution of timescales due to heterogeneity; whether this is connected to the dynamical heterogeneity in the entanglement dynamics, akin to classical glasses [76], remains a work for the future. Relatedly, the spacetime picture also raises the question that whether there exists a surface growth model for entanglement growth [16] which describes the dynamics of entanglement in the MBL phase.
@@ Acknowledgements.
We thank F. Alet, C. Artiaco, J. H. Bardarson, D. A. Chávez, N. Laflorencie and I. M. Khaymovich for several useful discussions. SR acknowledges support from SERB-DST, Government. of India under Grant No. SRG/2023/000858, from the Department of Atomic Energy, Government of India, under Project No. RTI4001, and from an ICTS-Simons Early Career Faculty Fellowship via a grant from the Simons Foundation (Grant No. 677895, R.G.).
@@ Appendix B Evidence for expression for $\tilde{\mathcal{P}}_{{\rm dyn}}$
Appendix C Derivation of the $\omega$-dependence of $\tilde{\mathcal{P}}^{AB}(\omega)$

### v2

@@ I Introduction
The question of if, and how, out-of-equilibrium quantum many-body systems thermalise as quantum correlations develop dynamically between distant degrees of freedom is one of the most fundamental questions of interest in condensed matter and statistical physics [1, 2, 3, 4, 5, 6]. This question is intimately connected to the entanglement structure and its dynamics [7, 8, 9] as it is one of the most fundamental manifestations of how quantum correlations and information propagates through the system [10, 11, 12, 13, 14, 15, 16, 17, 18]. Developing a microscopic theory for the dynamics of entanglement in quantum many-body systems is therefore naturally important.
@@ III Disordered Floquet spin-1/2 chain as a model
While the entire framework relating the dynamics of purity to eigenstate correlations described in the previous section was completely general, our main focus on strongly disordered systems in the MBL regime. In order to orient ourselves in that direction, and also as a concrete setting for demonstrating our theory, we employ a disordered, Floquet spin-1/2 chain. The Floquet unitary, $U_{F}$, is given by [67]
@@ IV.1 Special quartets which dominate dynamics of purity
The picture that therefore emerges for the structure of quartets of eigenstates with $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$ is as follows. The four eigenstates in the quartet are such that there localisation centres are made of two distinct $\sigma^{z}$-configurations in subsystem $A$ and two in $B$. In other words, the four distinct combinations of these states form the localisation centres of the four eigenstates. In fact, this picture straightforwardly explains the number of quartets with $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$. For a given |$\ket{\alpha}$, the number of states |$\ket{\beta}$ which has the same localisation centre in $B$ but different in $A$ is $N_{{\cal H}_{A}}-1$. Similarly, the number of states |$\ket{\gamma}$ which has the same localisation centre as |$\ket{\alpha}$ in $A$ but different in $B$ is $N_{{\cal H}_{B}}-1$. Fixing |$\ket{\beta}$ and |$\ket{\gamma}$ automatically fixes the state |$\ket{\lambda}$, see Eq. 28 or Fig. 5. Since there are $N_{\mathcal{H}}$ choices of |$\ket{\alpha}$ itself, the total number of quartets which satisfy Eq. 28, which we denote by ${\cal N}_{\square}$ is therefore $N_{\mathcal{H}}(N_{{\cal H}_{A}}-1)(N_{{\cal H}_{B}}-1)$. As the subsystems $A$ and $B$ correspond to left and right halves of the entire system, we have ${\cal N}_{\square}=N_{\mathcal{H}}(\sqrt{N_{\mathcal{H}}}-1)^{2}\approx N_{\mathcal{H}}^{2}$ or equivalently ${\cal N}_{\square}/{\cal N}_{Q}\sim N_{\mathcal{H}}^{-2}$. This immediately explains why a fraction $\sim N_{\mathcal{H}}^{-2}$ of the quartets have $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$ (see Fig. 4). It is worth clarifying and reiterating a couple of points here. First, our focus is deep in the MBL phase where associating an eigenstate to a localisation centre is well-defined despite the multifractality of the eigenstate. Second, the theoretical picture that emerges in the following sections does not depend on an eigenstate being sharply localised in the Hilbert space. Instead, the identification should be thought of as a starting point for constructing the theory [^2] In fact, the existence and the structure of the special quartets are generic to MBL systems; while the results in Fig. 4 and Fig. 5 are for the Floquet system in Eq. 23, an identical structure emerges also for time-independent MBL Hamiltonians as shown in Appendix C..
@@ IV.3 Emergence of logarithmic growth of entanglement
While the summations in Eq. 42 can be done exactly, as detailed in Appendix D, here we only sketch the derivation and discuss the physical import of the result. A couple of key assertions that we make here are $\mu>1$ and $\xi<1/\ln 2$; the numerical results in Fig. 10 indeed show that they are satisfied. However, we will see post facto that violating these bounds leads to results which are unphysical or incompatible with locality.
In the first summation in Eq. 42, note that the summand, $\sim 2^{-r}e^{r/\xi}$ is an exponentially increasing function of $r$ as $\xi<1/\ln 2$, and hence the sum is dominated by $r=r_{\ast}(\omega)$. As such, the first sum can be estimated as $2^{-r_{\ast}(\omega)}e^{r_{\ast}(\omega)/\xi}r_{\ast}(\omega)$ which using Eq. 41 is $\sim\omega^{-1+\xi\ln 2}\ln|\omega|$. On the other hand, in the second summation, by virtue of $\mu>1$ and $\xi<1/\ln 2$, the summands are an exponentially decreasing function of $r$, and hence the summation is expected to be dominated by the first term $r=r_{\ast}(\omega)+1$, which again leads to an estimate of the sum of as $\sim\omega^{-1+\xi\ln 2}|\ln(\omega/c)|$, which is the same as the first summation. We therefore conclude that
$\langle\tilde{\mathcal{P}}^{AB}(\omega)\rangle\sim\omega^{-1+\xi\ln 2}|\ln(\omega/c)|\,,$ (43)
@@
$\langle\mathcal{P}^{AB}(t)\rangle\sim t^{-\xi\ln 2}\ln(ct)\,,$ (44)
@@
$\langle S_{2}^{AB}(t)\rangle\sim(\xi\ln 2)\ln t-\ln\ln(ct)\overset{t\gg 1}{\sim}(\xi\ln 2)\ln t\,,$ (45)
which is the logarithmic growth of the entanglement entropy characteristic of the MBL regime. From the data in Fig. 10, we have $\xi=0.32$ for $\Gamma=0.1$ and $\xi=0.44$ for $\Gamma=0.15$. For these values, the power-law-exponent, $\xi\ln 2$ in Eq. 44, takes on values approximately 0.22 and 0.31 for $\Gamma=0.1$ and $0.15$ respectively. The numerical results in Fig. 2 on the other hand estimate the exponent $a$ to be 0.17 and 0.24. We note that while the two estimates are reasonably close, $a\lessapprox\xi\ln 2$ which we attribute to the logarithmic correction in Eq. 44. In particular, taking into account the logarithmic correction while analysing the data in Fig. 2 does yield power-law exponents in excellent agreement with the theoretical prediction, see Appendix E for details.
@@ IV.4 Spacetime picture of entanglement growth
An important physical import of the above analysis is that it provides a spacetime picture of the entanglement growth as follows. For a given $\omega$, the sum in Eq. 39 is dominated by a single $r=r_{\ast}(\omega)$. In other words, the dynamics of the purity at frequency $\omega$ is encoded in the spectral correlations of the form in Eq. 30 within the special quartets $\square_{r_{\ast}(\omega)}$ with $r_{\ast}(\omega)$ given by Eq. 41. This suggests that that at any time $t$, there is a characteristic lengthscale $d(t)\equiv r_{\ast}(\omega=2\pi/t)\sim\xi\ln t$ which contributes dominantly to the growth of entanglement at time $t$. Since degrees of freedom further away from the bipartition get entangled at later times, loosely speaking, $d(t)$ can be understood as a lengthscale degrees of freedom within which are entangled at time $t$. As the total entanglement carried in such a case will be proportional to $d(t)$, this presents an alternative perspective to infer the logarithmic growth of entanglement entropy. In addition, this could also be understood as the manifestation of the logarithmic light-cone of operator spreading in the MBL regime [44, 45, 46, 47].
Figure 11: The spacetime picture of entanglement growth obtained numerically. The heatmap shows $N_{r}{\cal P}_{r}(t)$ (defined in Eq. 46) in the $(r,t)$ plane. Note that the $t$ axis is logarithmic. As such the linear front seen in the plots corresponds to a logarithmic front in time. Data is for $L=12$ and the left and right panels correspond to $\Gamma=0.1$ and $0.15$ respectively. Also note that the wavefront is slower for smaller $\Gamma$ as one expects deeper in the MBL phase.
@@
where ${\cal P}_{r}(t)$ is the Fourier transform of $\tilde{\mathcal{P}}_{\square_{r}}(\omega)$. We first present microscopic results for $N_{r}{\cal P}_{r}(t)$ as a heatmap in $r$ and $t$, obtained directly from their numerical computation. The results are shown in Fig. 11 where it is clear that there is a ‘feature’ which spreads in space logarithmically in time (note that the time axes in the plots in Fig. 11 are logarithmic). In particular, it is clear that at a given $t$ there is a peak in $N_{r}{\cal P}_{r}(t)$ at a specific value of $r$ which grows logarithmic with $t$; this is nothing but the aforementioned $d(t)$. This provides us with the notion of an ‘entanglement front’.
To understand analytically the origins of this, note that the functional form of $\tilde{\mathcal{P}}_{\square_{r}}(\omega)$ in Eq. 40 suggests that ${\cal P}_{r}(t)$ decays like a stretched-exponential in time [71],
@@
An analytic expression of the kind above along with $N_{r}$ given by Eq. 35 allows us to show the entanglement wavefront in a much cleaner fashion for arbitrarily large systems, simply by plotting the closed form expressions. This is done in Fig. 12. Since $t_{\ast}(r)$ grows exponentially with $r$, the temporal dependence of ${\cal P}_{r}(t)$ naturally slows down with increasing $r$. At the same time, the number of quartets with $r$ also decays exponentially with $r$, Eq. 35. The power-law decay of $\langle\mathcal{P}^{AB}(t)\rangle$ emerges from the collective effect of the stretched-exponential decays in Eq. 47 with hierarchically growing timescales, $t_{\ast}(r)\sim e^{r/\xi}$ and the hierarchically decreasing $N_{r}\sim e^{-r\ln 2}$. This arises simply from the fact that the sum in Eq. 46 is dominated by a ‘saddle-point’, $r_{\ast}(t)\sim\xi\ln t$. It is also interesting to note that the phenomenology is quite similar to that of hierarchically constrained dynamics for glassy relaxation [72].
From the space-resolved picture discussed above, the entanglement wavefront can be understood as follows. The partial sum of Eq. 46, $\sum_{r=2}^{R}N_{r}{\cal P}_{r}(t)$ can be understood to contain the information of the decay of bipartite purity between $A$ and $B$ carried by degrees of freedom within a distance $2R$ from the bipartition. $N_{r}{\cal P}_{r}(t)$ is therefore the contribution to the decay in purity at time $t$ at distance $r$ from boundary. However, the aforementioned $r_{\ast}(t)$ suggests that there is a well-defined wavefront of entanglement spreading,as seen numerically in Fig. 11 and theoretically in Fig. 12. For a given $t$, the wavefront decays exponentially with $r$ away from $r_{\ast}(t)$ for $r>r_{\ast}(t)$ due to the exponential decay with $r$ of $N_{r}$. On the other hand, the hole in the profile for $r$ away from $r_{\ast}(t)$ with $r<r_{\ast}(t)$ is purely due to the fact that degrees of freedom within distance $r$ of the bipartition are strongly entangled and the effective contribution to the purity from them has decayed to extremely small values.
Figure 12: The spacetime picture of entanglement growth obtained from our theory. The left panel shows a heatmap in spacetime of $N_{r}{\cal P}_{r}(t)$, defined in Eq. 46, where we used the expression in Eq. 35 for $N_{r}$ and that in Eq. 47 for ${\cal P}_{r}(t)$. The wavefront spreads logarithmically in time as indicated by the red dashed line. The right panel shows wavefronts in space at fixed times (darker colours show later times) which are horizontal cuts of the figure in the left panel.
@@ VI Summary and Outlook
All the results summarised above are rooted in an understanding of the eigenstate and spectral correlations involving four different eigenstates. Questions to follow up in the future therefore arise where such correlations or their generalisations appear naturally. While we focused here on the entanglement entropy of a system with no conservation laws, a similar theory can be constructed for the unresolved issue of the dynamics of number entropy in MBL systems with a conserved charge [73, 74, 75, 76]. In order to better understand the spatiotemporal structure of statistics of eigenstates, analogous to Ref. [39], one can study the correlation $V_{\alpha\beta\gamma\lambda}(V_{\alpha\beta\gamma\lambda}^{\prime})^{\ast}$ resolved in $\theta_{\alpha\beta\gamma\lambda}$ where $V_{\alpha\beta\gamma\lambda}^{\prime}$ is the same eigenstate correlation as defined in Eq. 8 but with a different bipartition of the system. In a similar spirit, one can introduce matrix elements of local operators and recast the OTOC or temporal fluctuations of local observables [77] from the point-of-view of eigenstate correlations. In the same direction of spatiotemporal structure of entanglement dynamics, it will be important to understand the origins of the putative stretched-exponential decay in time of purity resolved in $r$ (see Eq. 47). Usually such anomalous decays appear due to the presence of a broad distribution of timescales due to heterogeneity; whether this is connected to the dynamical heterogeneity in the entanglement dynamics, akin to classical glasses [78], remains a work for the future. Relatedly, the spacetime picture also raises the question that whether there exists a surface growth model for entanglement growth [17] which describes the dynamics of entanglement in the MBL phase.
@@ Acknowledgements.
We thank F. Alet, C. Artiaco, J. H. Bardarson, D. A. Chávez, N. Laflorencie and I. M. Khaymovich for several useful discussions. SR acknowledges support from SERB-DST, Government of India under Grant No. SRG/2023/000858, from the Department of Atomic Energy, Government of India, under Project No. RTI4001, and from an ICTS-Simons Early Career Faculty Fellowship via a grant from the Simons Foundation (Grant No. 677895, R.G.).
@@ Appendix B Evidence for expression for $\tilde{\mathcal{P}}_{{\rm dyn}}$
Appendix C Existence and structure of special quartets for time-independent MBL Hamiltonians
As discussed in Sec. IV, the central ingredient of our theory the existence dominant quartets whose multiplicity scales as $N_{\mathcal{H}}^{2}$ and for which $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$. While the results centred on the Floquet system defined in Eq. 23, here we show that they hold also for a time-independent MBL Hamiltonian, thereby providing evidence that the existence and structure of the dominant quartets is genetic to MBL phases.
We consider the Hamiltonian [79],
$H=\sum_{i}[J_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}+h_{i}\sigma^{z}_{i}+\Gamma\sigma^{x}_{i}]\,,$ (66)
where $J_{i}\in[0.8,1.2]$, $\Gamma=1$, and $h_{i}\in[-W,W]$ We choose $W=10.0$ to be sufficiently deep in the MBL phase and present plots analogous to those in Fig. 4 and Fig. 5. The results for the Hamiltonian (66) are shown in Fig. 14 and Fig. 15. The former is identical to Fig. 4 whereas the latter is identical to Fig. 5, thus showing that the multiplicity and the structure of the dominants quartets for the Hamiltonian (66) are identical to that of the Floquet system (23).
Figure 14: Data analogous to Fig. 4 for the time-independent Hamiltonian in Eq. 66, again showing that there exist dominant quartets with $|V_{\alpha\beta\gamma\lambda}|^{2}\sim O(1)$ whose multiplicity scales as $N_{\mathcal{H}}^{2}$.
Figure 15: Data analogous to that of Fig. 5, but for the Hamiltonian in Eq. 66, again shows that the structure of the dominants quartets in both, Fock space and real space is the same.
Appendix D Derivation of the $\omega$-dependence of $\tilde{\mathcal{P}}^{AB}(\omega)$